# Universal quantum control over bosonic network

### v1

@@ III State exchange between two bosonic modes
Similarly, in Fig. 1(c), the mode $a_{1}$ is prepared as a cat state $|\rm cat\rangle=(|\alpha\rangle+|-\alpha\rangle)/\mathcal{N}$ with $\alpha=5$ and $\mathcal{N}$ the normalization coefficient. Again it is confirmed that $a_{2}(\tau)=a_{1}(0)$ and $a_{1}(\tau)=a_{2}(0)$. Our protocol even applies to the mixed state. In Fig. 1(d), the fidelity is evaluated by $F={\rm Tr}[\rho(t)\rho_{\rm th}]$, where $\rho(t)$ is the solution to the von-Neumann equation driven by the original Hamiltonian (19) and $\rho_{\rm th}=\sum_{n}p_{n}|n\rangle\langle n|$ with $p_{n}=(\bar{n}^{n})/(1+\bar{n})^{n+1}$ with $\bar{n}=1$ the average occupation. A complete exchange is also observed for the thermal states $\rho_{\rm th}$ and $|0\rangle\langle 0|$, as shown in Fig. 1(d).
IV Chiral state transfer among three modes
@@
where the coupling strengths are $J_{1}$, $J_{2}$, and $J_{3}$ and the phases are $\varphi_{1}$, $\varphi_{2}$, and $\varphi_{3}$, respectively. In the rotating frame with respect to $H_{0}=\omega_{0}(t)/2(a_{1}^{\dagger}a_{1}+a_{2}^{\dagger}a_{2}+a_{3}^{\dagger}a_{3})$, we have
@@
$V_{\alpha_{2}}(t)$ $=e^{-i\frac{\delta\alpha_{2}}{2}\left[b_{1}^{\dagger}(0)b_{1}(0)-a_{3}^{\dagger}a_{3}\right]},$ (35)
$V_{\theta_{2}}(t)$ $=e^{-\delta\theta_{2}[e^{i\beta(0)}a_{3}^{\dagger}b_{1}(0)-e^{-i\beta(0)}b_{1}^{\dagger}(0)a_{3}]}$
with $\delta\alpha_{2}=\alpha_{2}(t)-\alpha_{2}(0)$ and $\delta\theta_{2}=\theta_{2}(t)-\theta_{2}(0)$.
@@
Under the conditions in Eqs. (36), (37) and (38), the ancillary modes $\mu_{k}(t)$’s in Eq. (32) can be activated as useful nonadiabatic passages. The ancillary modes $\mu_{k}(0)$ evolve with time in accordance to Eq. (18) as
@@
During the practical control, one can chose $\theta_{1}(t)$, $\theta_{2}(t)$, $f_{1}(t)$, and $f(t)$ as independent variables to avoid the singularity of the experimental parameters. Specifically, using Eq. (41), Eq. (38) can be rewritten as
@@
In particular, during Stage (i), under the boundary conditions of $\theta_{1}(0)=0$, $\theta_{2}(0)=0$, $\theta_{1}(\tau)=\pi/2$ and $\theta_{2}(\tau)=\pi/2$, we have $\mu_{1}(0)=a_{1}\rightarrow\mu_{1}(\tau)=a_{2}$, $\mu_{2}(0)=a_{2}\rightarrow\mu_{2}(\tau)=a_{3}$, and $\mu_{3}(0)=a_{3}\rightarrow\mu_{1}(\tau)=a_{1}$ according to Eq. (33). It indicates that the states of modes $a_{1}$, $a_{2}$, and $a_{3}$ at $t=0$, are transferred to $a_{2}$, $a_{3}$, and $a_{1}$, respectively, at $t=\tau$. Then the bases in $|\psi(0)\rangle$ are transformed as
@@
$\theta_{1}(t)$ $=\frac{\pi(t-k\tau)}{2\tau},\quad\theta_{2}(t)=\theta_{1}(t),$ (46)
$\theta_{1}(t)$ $=\frac{\pi(t-k\tau)}{2\tau}-\frac{\pi}{2},\quad\theta_{2}(t)=\theta_{1}(t),$
$\theta_{1}(t)$ $=\frac{\pi(t-k\tau)}{2\tau}+\pi,\quad\theta_{2}(t)=\theta_{1}(t)-\frac{\pi}{2}$
@@
$\theta_{1}(t)$ $=\frac{\pi(t-k\tau)}{2\tau}+\frac{\pi}{2},\quad\theta_{2}(t)=\theta_{1}(t)-\frac{\pi}{2},$ (47)
$\theta_{1}(t)$ $=\frac{\pi(t-k\tau)}{2\tau},\quad\theta_{2}(t)=\theta_{1}(t)+\frac{\pi}{2},$
$\theta_{1}(t)$ $=\frac{\pi(t-k\tau)}{2\tau}+\frac{\pi}{2},\quad\theta_{2}(t)=\theta_{1}(t)+\frac{\pi}{2},$
@@
The performance of our protocol can be evaluated by the dynamics of the state or entanglement fidelity $\mathcal{F}_{jk}(t)=|\langle\psi(t)|\phi(2)\rangle_{jk}|0\rangle_{l\neq j,k}|^{2}$ with respect to the target state $|\phi(2)\rangle_{jk}$ defined in Eq. (44), where $|\psi(t)\rangle$ is the pure-state solution of the Schrödinger equation with the original Hamiltonian (30). In Fig. 3(a), it is found that the perfect counterclockwise entangled state transfer can be achieved as $\mathcal{F}_{13}(0)=1$ when $t=0$, $\mathcal{F}_{12}(\tau)=1$ when $t=\tau$, $\mathcal{F}_{23}(2\tau)=1$ when $t=2\tau$ and $\mathcal{F}_{13}(3\tau)=1$ when $t=3\tau$. During the period $t\in[3\tau,6\tau]$, the second loop perfectly repeats the first one. In Fig. 3(b), the NOON state is perfectly transferred in a clockwise manner. Specifically, we have $\mathcal{F}_{13}(0)=1$ when $t=0$, $\mathcal{F}_{23}(\tau)=1$ when $t=\tau$, $\mathcal{F}_{12}(2\tau)=1$ when $t=2\tau$, and $\mathcal{F}_{13}(3\tau)=1$ when $t=3\tau$. The behavior of the second loop is also identical to that of the first loop. Note the chiral transfer of continuous-variable system in the current protocol is realized through nonadiabatic control rather than the Floquet driving Qi and Jing (2022b) that is featured with a fixed ratio of the driving intensity and frequency and the uniformly distributed local phases. Thus, the nonadiabatic passage is much flexible in parametric setting than the Floquet driving, e.g., Eqs. (46) and (47) can be replaced with any functions following the same boundary conditions. Also, the current protocol enables the chiral transfer of selected nodes in the whole network, without eliminating the unwanted couplings or connections.
@@ V Chiral state transfer in network
In this section, our control protocol is applied to a central-configuration bosonic network as shown in Fig. 4 which consists of $N$ bosonic modes. The central node $a_{N}$ is coupled to the other modes $a_{n}$’s, $1\leq n\leq N-1$, via the exchange interaction that is characterized by the coupling strength $J_{n}$ and the phase $\varphi_{n}$. Then the full Hamiltonian can be written as
@@
Here, to clarify the underlying ideas of our theory, we would further elaborate how the framework introduced in Sec. II can be applied to any $N$-mode system, such as the centralized network in Fig. 4 or the bosonic network composed of two coupled subsystems Huang et al. (2023) if there exist more than one central node. In our framework, the dynamics of the general system can be described in the time-dependent ancillary modes $\mu_{k}(t)$ in Eq. (3). By applying the unitary transformation $V_{N-1}(t)$ in Eq. (8) together with the commutation condition in Eq. (13), the resulting constraints on the Hamiltonian (49) activate the ancillary modes in Eq. (18) to be universal passages for versatile tasks, e.g., exchanging arbitrary unknown states between the desired bosonic modes in a fully connected network, irrespective of the network size.
As a concrete example, we consider the Hamiltonian with $N=4$ in Eq. (49). Then the ancillary modes in Eq. (3) is written as
@@
$V_{3}(t)=V_{\alpha_{1}}V_{\theta_{1}}V_{\alpha_{2}}V_{\theta_{2}}V_{\alpha_{3}}V_{\theta_{3}}=\prod_{1}^{k=3}V_{\alpha_{k}}V_{\theta_{k}}$ (51)
@@
Along the passage $\mu_{4}(t)$ with the parameters subject to Eqs. (52) and (LABEL:CondDetu). A chiral state transfer among the modes $a_{1}$, $a_{2}$, and $a_{3}$ can be divided into three stages of state conversion, e.g., in Stage (i) for $t\in[0,\tau]$, $a_{1}\leftrightarrow a_{2}$, in Stage (ii) for $t\in[\tau,2\tau]$, $a_{2}\leftrightarrow a_{3}$, and in Stage (iii) for $t\in[2\tau,3\tau]$, $a_{3}\leftrightarrow a_{1}$. In other words, an arbitrary initial state in mode $a_{1}$ propagates to mode $a_{2}$ at $t=\tau$, to mode $a_{3}$ at $t=2\tau$, and back to mode $a_{1}$ at $t=3\tau$, despite they are not directly connecting with each other. This task requires the parameters $\theta_{n}(t)$’s with $1\leq n\leq 3$ to satisfy the boundary conditions: $\theta_{1}(0)=\theta_{2}(0)=\theta_{3}(0)=\pi/2$, $\theta_{1}(\tau)=\pi$, $\theta_{2}(\tau)=\theta_{3}(\tau)=\pi/2$, $\theta_{2}(2\tau)=0$, $\theta_{3}(2\tau)=\pi/2$, and $\theta_{1}(3\tau)=\theta_{2}(3\tau)=\theta_{3}(3\tau)=\pi/2$. Then one can check that $\mu_{4}(0)=a_{1}\rightarrow\mu_{4}(\tau)=a_{2}\rightarrow\mu_{4}(2\tau)=a_{3}\rightarrow\mu_{4}(3\tau)=a_{1}$. In general, $\theta_{1}(t)$ and $\theta_{2}(t)$ during Stages (i), (ii), and (iii) of the $k$th loop, $k\geq 1$, can be set as
$\theta_{1}(t)$ $=\theta_{2}(t)+\frac{\pi}{2},\quad\theta_{2}(t)=2\Phi(t)+\frac{\pi}{2},$ (54)
@@
Figure 5: Population dynamics $P_{n}(t)$ for (a) partial states with five excitations in the state transfer $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|5\rangle_{2}|0\rangle_{3}|0\rangle_{4}$ and (b) the target states in chiral population transfer $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|5\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|0\rangle_{2}|5\rangle_{3}|0\rangle_{4}\rightarrow|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$. The coupling strengths and the detunings are set as Eqs. (52) and (LABEL:CondDetu) under $\varphi_{1}+\alpha_{3}=\pi/2$, $\varphi_{2}-\alpha_{1}+\alpha_{3}=\pi/2$, $\varphi_{3}-\alpha_{2}+\alpha_{3}=\pi/2$, $\theta_{1}(t)$ and $\theta_{2}(t)$ in Eq. (54), and $\theta_{3}(t)$ in Eq. (55).
@@
For example, we consider the chiral propagation of the Fock state $|n\rangle$ among the three modes. Assuming that the whole system is initially in the state $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$, i.e., the mode $a_{1}$ is prepared in $|n\rangle$ with $n=5$ and the other modes are in the vacuum state $|0\rangle$. Our protocol can be evaluated by the population $P_{n}=|\langle n|\psi(t)\rangle|^{2}$ where $|n\rangle$ denotes some of the relevant number states with conserved excitations and $|\psi(t)\rangle$ is obtained by numerical simulation. Figure 5(a) presents the population dynamics $P_{n}(t)$ about the state conversion between the modes $a_{1}$ and $a_{2}$ in Stage (i). It is found that the initial population in the state $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$ can be completely transferred to the state $|0\rangle_{1}|5\rangle_{2}|0\rangle_{3}|0\rangle_{4}$ when $t=\tau$, despite the other intermediate states can be temporally occupied during Stage (i). For instance, when $t=0.5\tau$, the populations are $P_{0023}(0.5\tau)=0.311$, $P_{0032}(0.5\tau)=0.314$, $P_{0014}(0.5\tau)=0.154$, $P_{0041}(0.5\tau)=0.158$, and $P_{0050}=0.032$. Two completed loops of chiral Fock-state transfer, i.e., $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|5\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|0\rangle_{2}|5\rangle_{3}|0\rangle_{4}\rightarrow|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$ are presented in Fig. 5(b). We find that the perfect transfer can be realized as $P_{5000}(0)=1$, $P_{0500}(\tau)=1$, $P_{0050}(2\tau)=1$, and $P_{5000}(3\tau)=1$. The second loop behaves the same as the first one.

### v2

@@ III State exchange between two bosonic modes
Similarly, in Fig. 1(c), the mode $a_{1}$ is initially prepared as a cat state $|\rm cat\rangle=(|\alpha\rangle+|-\alpha\rangle)/\mathcal{N}$ with $\alpha=5$ and $\mathcal{N}$ the normalization coefficient. Again it is confirmed that $a_{2}(\tau)=a_{1}(0)$ and $a_{1}(\tau)=a_{2}(0)$. Our protocol even applies to the mixed state. In Fig. 1(d), the fidelity is evaluated by $F={\rm Tr}[\rho(t)\rho_{\rm th}]$, where $\rho(t)$ is the solution to the von-Neumann equation driven by the original Hamiltonian (19) and $\rho_{\rm th}=\sum_{n}p_{n}|n\rangle\langle n|$ with $p_{n}=(\bar{n}^{n})/(1+\bar{n})^{n+1}$ with $\bar{n}=1$ the average occupation. A complete exchange is also observed for the thermal states $\rho_{\rm th}$ and $|0\rangle\langle 0|$, as shown in Fig. 1(d).
Figure 2: Fidelity dynamics $F(t)$ for the target state $|0,\rm cat\rangle$ during the exchange of cat state $|\rm cat,0\rangle\rightarrow|0,\rm cat\rangle$: (a) for various environmental decay rates $\kappa$; and (b) for various derivation coefficients $\epsilon$ in the imperfect parameter setting $\theta_{1}(t)\rightarrow(1+\epsilon)\theta_{1}(t)$ and $\theta_{1}(t)=(\pi t)/(2\tau)$. The evolution period is set as $\tau=100$ ns and the coupling strength $J/2\pi\sim 25$ MHz. The other parameters are the same as Fig. 1(c).
The robustness of our protocol can be discussed against the environmental coupling or the parameter imperfection. The system dynamics in the presence of decoherence can be described by the master equation Carmichael (1999),
$\frac{\partial}{\partial t}\rho=-i[H(t),\rho]+\frac{\kappa}{2}\mathcal{L}_{a_{1}}[\rho]+\frac{\kappa}{2}\mathcal{L}_{a_{2}}[\rho],$ (30)
where $H(t)$ is the ideal Hamiltonian in Eq. (19) and the Lindblad superoperators are defined as $\mathcal{L}_{o}[\rho]=2o\rho o^{\dagger}-o^{\dagger}o\rho-\rho o^{\dagger}o$ with $o=a_{1},a_{2}$ indicating the decay channels for the bosonic modes $a_{1}$ and $a_{2}$, respectively. The decay rates are set as the same value. In the state-of-art cavity magnonic systems Lachance-Quirion et al. (2019); Zare Rameshti et al. (2022), the exchange coupling strength between the cavity and magnon modes can be as large as $J/2\pi\sim 25$ MHz Lachance-Quirion et al. (2019); Zare Rameshti et al. (2022) and the decay rate is in the order of $\kappa/2\pi\sim 1$ MHz. In Fig. 2(a), we demonstrate the fidelity dynamics of the target state $|0,\rm cat\rangle$ during the cat state exchange of $|\rm cat,0\rangle\rightarrow|0,\rm cat\rangle$ between the bosonic modes $a_{1}$ and $a_{2}$. Our protocol is found to be robust against decoherence. In particular, we have $F_{0,\rm cat}(\tau)=0.993$ when $\kappa\tau/2\pi\sim 0.05$ (close to the experimental conditions Lachance-Quirion et al. (2019); Zare Rameshti et al. (2022)) and $F_{0,\rm cat}(\tau)=0.867$ when $\kappa\tau/2\pi\sim 0.1$. In Fig. 2(b), we plot the same fidelity under the parametric fluctuation, i.e., $\theta_{1}(t)\rightarrow(1+\epsilon)\theta_{1}(t)$ with the deviation coefficient $\epsilon$, which affects both detuning and coupling strength due to Eqs. (28a) and (28c). It is found that our protocol is also insensitive to the imperfect parameter setting. In particular, we have $F_{0,\rm cat}(\tau)=0.999$ when $\epsilon=0.01$, $F_{0,\rm cat}(\tau)=0.982$ when $\epsilon=0.05$, and $F_{0,\rm cat}(\tau)=0.929$ when $\epsilon=0.1$.
IV Chiral entanglement transfer among three modes
@@
where $J_{j}$ and $\varphi_{j}$ are coupling strength and phase, respectively. In the rotating frame with respect to $H_{0}=\omega_{0}(t)/2(a_{1}^{\dagger}a_{1}+a_{2}^{\dagger}a_{2}+a_{3}^{\dagger}a_{3})$, we have
@@
$V_{\alpha_{2}}(t)$ $=e^{-i\frac{\alpha_{2}(t)-\alpha_{2}(0)}{2}\left[b_{1}^{\dagger}(0)b_{1}(0)-a_{3}^{\dagger}a_{3}\right]},$ (36)
$V_{\theta_{2}}(t)$ $=e^{-[\theta_{2}(t)-\theta_{2}(0)][e^{i\alpha_{2}(0)}a_{3}^{\dagger}b_{1}(0)-e^{-i\alpha_{2}(0)}b_{1}^{\dagger}(0)a_{3}]}.$
@@
Under the constraint conditions in Eqs. (37), (38), and (39), the ancillary modes $\mu_{k}(t)$’s in Eq. (33) can be activated as useful nonadiabatic passages. The ancillary modes $\mu_{k}(0)$’s evolve with time in accordance to Eq. (18) as
@@
During the practical control, one can choose $\theta_{1}(t)$, $\theta_{2}(t)$, $f_{1}(t)$, and $f(t)$ as independent variables to avoid the singularity of the experimental parameters. Specifically, using Eq. (42), Eq. (39) can be rewritten as
@@
In particular, during Stage (i), under the boundary conditions of $\theta_{1}(0)=0$, $\theta_{2}(0)=0$, $\theta_{1}(\tau)=\pi/2$, and $\theta_{2}(\tau)=\pi/2$, we have $\mu_{1}(0)=a_{1}\rightarrow\mu_{1}(\tau)=a_{2}$, $\mu_{2}(0)=a_{2}\rightarrow\mu_{2}(\tau)=a_{3}$, and $\mu_{3}(0)=a_{3}\rightarrow\mu_{1}(\tau)=a_{1}$ according to Eq. (33). It indicates that the states of modes $a_{1}$, $a_{2}$, and $a_{3}$ at $t=0$, are transferred to $a_{2}$, $a_{3}$, and $a_{1}$, respectively, at $t=\tau$. Then the bases in $|\psi(0)\rangle$ are transformed as
@@
$\theta_{1}(t)$ $=\frac{\pi[t-3(k-1)\tau]}{2\tau},\quad\theta_{2}(t)=\theta_{1}(t),$ (47)
$\theta_{1}(t)$ $=\frac{\pi[t-3(k-1)\tau]}{2\tau}-\frac{\pi}{2},\quad\theta_{2}(t)=\theta_{1}(t),$
$\theta_{1}(t)$ $=\frac{\pi[t-3(k-1)\tau]}{2\tau}+\pi,\quad\theta_{2}(t)=\theta_{1}(t)-\frac{\pi}{2}$
@@
$\theta_{1}(t)$ $=\frac{\pi[t-3(k-1)\tau]}{2\tau}+\frac{\pi}{2},\quad\theta_{2}(t)=\theta_{1}(t)-\frac{\pi}{2},$ (48)
$\theta_{1}(t)$ $=\frac{\pi[t-3(k-1)\tau]}{2\tau},\quad\theta_{2}(t)=\theta_{1}(t)+\frac{\pi}{2},$
$\theta_{1}(t)$ $=\frac{\pi[t-3(k-1)\tau]}{2\tau}+\frac{\pi}{2},\quad\theta_{2}(t)=\theta_{1}(t)+\frac{\pi}{2},$
@@
The performance of our protocol can be evaluated by the dynamics of the state or entanglement fidelity $\mathcal{F}_{jk}=|\langle\psi(t)|\phi(2)\rangle_{jk}|0\rangle_{l\neq j,k}|^{2}$ with respect to the target state $|\phi(2)\rangle_{jk}$ defined in Eq. (45), where $|\psi(t)\rangle$ is the pure-state solution of the Schrödinger equation with the original Hamiltonian (31). In Fig. 4(a), it is found that the perfect counterclockwise entanglement transfer can be achieved as $\mathcal{F}_{13}(0)=1$ when $t=0$, $\mathcal{F}_{12}(\tau)=1$ when $t=\tau$, $\mathcal{F}_{23}(2\tau)=1$ when $t=2\tau$ and $\mathcal{F}_{13}(3\tau)=1$ when $t=3\tau$. During the period $t\in[3\tau,6\tau]$, the second loop perfectly repeats the first one. In Fig. 4(b), the NOON state is perfectly transferred in a clockwise manner. Specifically, we have $\mathcal{F}_{13}(0)=1$ when $t=0$, $\mathcal{F}_{23}(\tau)=1$ when $t=\tau$, $\mathcal{F}_{12}(2\tau)=1$ when $t=2\tau$, and $\mathcal{F}_{13}(3\tau)=1$ when $t=3\tau$. The behavior of the second loop is also identical to that of the first loop. Note the chiral transfer of continuous-variable system in the current protocol is realized through nonadiabatic control rather than the Floquet driving Qi and Jing (2022b) that was featured with a fixed ratio of the driving intensity and frequency and the uniformly distributed local phases. Thus, the nonadiabatic passage is much flexible in parametric setting than the Floquet driving, e.g., Eqs. (47) and (48) can be replaced with any functions holding the same boundary conditions. Also, the current protocol enables the chiral transfer of selected nodes in the entire network, without eliminating the irrelevant couplings or connections.
@@ V Chiral state transfer in network
In this section, our control protocol is applied to a central-configuration bosonic network as shown in Fig. 5 which consists of $N$ bosonic modes. The central node $a_{N}$ is coupled to the other $a_{n}$’s, $1\leq n\leq N-1$, via the exchange interaction that is characterized by the coupling strength $J_{n}$ and the phase $\varphi_{n}$. Then the full Hamiltonian can be written as
@@
Here, to clarify the underlying ideas of our theory, we would further elaborate how the framework introduced in Sec. II can be applied to any $N$-mode system, such as the centralized network in Fig. 5 or the bosonic network composed of two coupled subsystems Huang et al. (2023) if there exist more than one central node. In our framework, the dynamics of the general system can be described in the time-dependent ancillary modes $\mu_{k}(t)$ in Eq. (3). By applying the unitary transformation $V_{N-1}(t)$ in Eq. (8) together with the commutation condition in Eq. (13), the resulting constraints on the Hamiltonian (50) activate the ancillary modes in Eq. (18) to be universal passages for versatile tasks, e.g., exchanging arbitrary unknown states between the desired bosonic modes in a connected network, irrespective of the network size.
As a concrete example, we consider the Hamiltonian in Eq. (50) with $N=4$. Then the ancillary modes in Eq. (3) is written as
@@
$V_{3}(t)=V_{\alpha_{1}}V_{\theta_{1}}V_{\alpha_{2}}V_{\theta_{2}}V_{\alpha_{3}}V_{\theta_{3}}=\prod_{k=1}^{3}V_{\alpha_{k}}V_{\theta_{k}}$ (52)
@@
Along the passage $\mu_{4}(t)$ with the constraint conditions in Eqs. (53) and (LABEL:CondDetu), a chiral state transfer among the modes $a_{1}$, $a_{2}$, and $a_{3}$ can be divided into three stages of state conversion, i.e., in Stage (i) for $t\in[0,\tau]$, $a_{1}\leftrightarrow a_{2}$, in Stage (ii) for $t\in[\tau,2\tau]$, $a_{2}\leftrightarrow a_{3}$, and in Stage (iii) for $t\in[2\tau,3\tau]$, $a_{3}\leftrightarrow a_{1}$. In other words, an arbitrary initial state in mode $a_{1}$ propagates to mode $a_{2}$ at $t=\tau$, to mode $a_{3}$ at $t=2\tau$, and back to mode $a_{1}$ at $t=3\tau$, despite they have no direction connections. This task requires the parameters $\theta_{n}(t)$’s with $1\leq n\leq 3$ to satisfy the boundary conditions: $\theta_{1}(0)=\theta_{2}(0)=\theta_{3}(0)=\pi/2$, $\theta_{1}(\tau)=\pi$, $\theta_{2}(\tau)=\theta_{3}(\tau)=\pi/2$, $\theta_{2}(2\tau)=0$, $\theta_{3}(2\tau)=\pi/2$, and $\theta_{1}(3\tau)=\theta_{2}(3\tau)=\theta_{3}(3\tau)=\pi/2$. Then one can check that $\mu_{4}(0)=a_{1}\rightarrow\mu_{4}(\tau)=a_{2}\rightarrow\mu_{4}(2\tau)=a_{3}\rightarrow\mu_{4}(3\tau)=a_{1}$. In general, $\theta_{1}(t)$ and $\theta_{2}(t)$ during Stages (i), (ii), and (iii) of the $k$th loop, $k\geq 1$, can be set as
$\theta_{1}(t)$ $=\Phi(t)+\frac{\pi}{2},\quad\theta_{2}(t)=2\Phi(t)+\frac{\pi}{2},$ (55)
@@
Figure 6: Population dynamics $P_{n}(t)$ for (a) partial states with five excitations in the state transfer $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|5\rangle_{2}|0\rangle_{3}|0\rangle_{4}$ and (b) the target states in chiral population transfer $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|5\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|0\rangle_{2}|5\rangle_{3}|0\rangle_{4}\rightarrow|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$. Coupling strengths and detunings are set as Eqs. (53) and (LABEL:CondDetu) under $\varphi_{1}+\alpha_{3}=\pi/2$, $\varphi_{2}-\alpha_{1}+\alpha_{3}=\pi/2$, $\varphi_{3}-\alpha_{2}+\alpha_{3}=\pi/2$, $\theta_{1}(t)$ and $\theta_{2}(t)$ in Eq. (55), and $\theta_{3}(t)$ in Eq. (56).
@@
We assume that the whole system is initially in the Fock state $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$, i.e., the mode $a_{1}$ is prepared in $|5\rangle$ and the other modes are in the vacuum state $|0\rangle$. Our protocol is evaluated by the population $P_{n}=|\langle n|\psi(t)\rangle|^{2}$, where $|n\rangle$’s are a group of relevant number states with conserved excitations and $|\psi(t)\rangle$ is obtained by numerical simulation. Figure 6(a) presents the dynamics about the population on $15$ states during Stage (i). It is found that the initial population on $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$ can be completely transferred to the state $|0\rangle_{1}|5\rangle_{2}|0\rangle_{3}|0\rangle_{4}$ when $t=\tau$, despite the other intermediate states can be temporally occupied. For instance, when $t=0.5\tau$, the populations are $P_{0023}(0.5\tau)=0.31$, $P_{0032}(0.5\tau)=0.31$, $P_{0014}(0.5\tau)=0.15$, $P_{0041}(0.5\tau)=0.16$, and $P_{0050}=0.032$. We plot two completed loops of chiral Fock-state transfer, i.e., $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|5\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|0\rangle_{2}|5\rangle_{3}|0\rangle_{4}\rightarrow|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$, in Fig. 6(b). It is found that the perfect transfer can be realized as $P_{5000}(0)=1$, $P_{0500}(\tau)=1$, $P_{0050}(2\tau)=1$, and $P_{5000}(3\tau)=1$. The second loop behaves the same as the first one.
Figure 7: Population dynamics $P_{n}(t)$ for the target states in chiral population transfer $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|5\rangle_{2}|0\rangle_{3}|0\rangle_{4}\rightarrow|0\rangle_{1}|0\rangle_{2}|5\rangle_{3}|0\rangle_{4}\rightarrow|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$ in the presence of unwanted coupling $J_{e}\sim\lambda{\rm max}\{J_{n}\}\times 10^{-2}$, $1\leq n\leq 3$ with (a) $\lambda\sim 0.5$, (b) $\lambda\sim 1$, and (c) $\lambda\sim 5$. In comparison, the result with a known $J_{e}$ is shown in (d). The coupling strengths $J_{n}$ with $1\leq n\leq 3$ and $J_{e}$ in (d) are constrained by Eqs. (58) and (59), while the other parameters in (a), (b), (c), and (d) are the same as Fig. 6(b).
In the presence of unwanted and unknown crosstalk among the network nodes, our protocol demonstrates a certain degree of robustness. For example, consider an extra coupling $J_{e}$ between the modes $a_{1}$ and $a_{2}$, the nonideal Hamiltonian can be written as the sum of the ideal Hamiltonian (50) with $N=4$ and the extra interaction Hamiltonian:
$H(t)\rightarrow H(t)+H_{1}(t),\quad H_{1}(t)=J_{e}\left(a_{1}^{\dagger}a_{2}+a_{2}^{\dagger}a_{1}\right).$ (57)
In the current cavity magnonic systems with multiple modes Zhang et al. (2015b), it is experimentally feasible to couple two magnon modes to a common cavity mode while keeping the direct or unwanted magnon-magnon coupling negligible. Therefore $J_{e}$ can be about two orders of magnitude lower than the intended coupling strength, i.e., $J_{e}=\lambda{\rm max}\{J_{n}\}\times 10^{-2}$, $1\leq n\leq 3$, with $\lambda$ a dimensionless factor measuring the unwanted connection strength. In Figs. 7(a), (b) and (c), we plot the population dynamics $P_{n}$ during the first loop about the chiral Fock-state transfer among the modes $a_{1}$, $a_{2}$, and $a_{3}$ with various $\lambda$. The network initial state remains as $|5\rangle_{1}|0\rangle_{2}|0\rangle_{3}|0\rangle_{4}$. We still have an almost perfect chiral transfer under a weak unwanted connection. In particular, in Fig. 7(a) with $\lambda=0.5$, we obtain $P_{0500}(\tau)=0.999$, $P_{0050}(2\tau)=0.997$, and $P_{5000}(3\tau)=0.995$. In Fig. 7(b) with $\lambda=1$, $P_{0500}(\tau)=0.957$, $P_{0050}(2\tau)=0.902$, and $P_{5000}(3\tau)=0.880$. When $\lambda$ is enhanced to $\lambda=5$ as shown in Fig. 7(c), the fidelities are found to be as low as $P_{0500}(\tau)=0.772$, $P_{0050}(2\tau)=0.576$, and $P_{5000}(3\tau)=0.571$.
Nevertheless, a perfect cyclic transfer can be still attainable by the evolution along the passage $\mu_{4}(t)$ if $J_{e}$ is known. In this case, the commutation condition (13) with the nonideal Hamiltonian (57) gives rise to the constraint equations deviated from Eqs. (53) and (LABEL:CondDetu). Specifically, we have $\Delta_{n}(t)=0$, $1\leq n\leq 4$, for the detunings, $\varphi_{1}=\varphi_{2}=\varphi_{3}=\pi/2$ for the phases, and Eq. (53) for coupling strengths are modified to be
$J_{1}$ $=J\sin\theta_{1},$ (58)
$J_{2}$ $=J\cos\theta_{1},$
$J_{3}$ $=-\dot{\theta}_{3}\cos\theta_{2}+\dot{\theta}_{2}\tan\theta_{3}\sin\theta_{2},$
$J$ $=-\dot{\theta}_{3}\sin\theta_{2}-\dot{\theta}_{2}\tan\theta_{3}\cos\theta_{2},$
where $J$ serves as a scaling coupling strength. The constraint equation for the extra coupling strength $J_{e}$ is found to be
$J_{e}=-\dot{\theta}_{1},$ (59)
which can be a constant as long as $\theta_{1}(t)$ is a linear function with time. Then, again with the boundary conditions given by Eqs. (55) and (56), any initial state in the mode $a_{1}$ can be perfectly transferred to the mode $a_{2}$ at $t=\tau$, to the mode $a_{3}$ at $t=2\tau$, and returns to the mode $a_{1}$ at $t=3\tau$, along the passage $\mu_{4}(t)$. The relevant population dynamics $P_{n}(t)$ with a known $J_{e}$ presents in Fig. 7(d) in comparison to Figs. 7(a), (b), and (c).